How many grams of copper are required to make a hollow shell with inner radius of 5.70cm and an outer radius of 5.75?The density of copper 8.93g/cm^3
1
Expert's answer
2020-06-22T11:08:36-0400
"m=\rho V=\rho\left(\frac{4}{3}\pi R^3-\frac{4}{3}\pi r^3\right)"

"m=8930\left(\frac{4}{3}\pi (0.0575)^3-\frac{4}{3}\pi (0.057)^3\right)=0.184\ kg"